### v1

# Watch Your Language: Large Language Models and Content Moderation

Abstract
Large language models (LLMs) have exploded in popularity due to their ability to perform a wide array of natural language tasks. Text-based content moderation is one LLM use case that has received recent enthusiasm, however, there is little research investigating how LLMs perform in content moderation settings. In this work, we evaluate a suite of modern, commercial LLMs (GPT-3, GPT-3.5, GPT-4) on two common content moderation tasks: rule-based community moderation and toxic content detection. For rule-based community moderation, we construct 95 LLM moderation-engines prompted with rules from 95 Reddit subcommunities and find that LLMs can be effective at rule-based moderation for many communities, achieving a median accuracy of 64% and a median precision of 83%. For toxicity detection, we find that LLMs significantly outperform existing commercially available toxicity classifiers. However, we also find that recent increases in model size add only marginal benefit to toxicity detection, suggesting a potential performance plateau for LLMs on toxicity detection tasks. We conclude by outlining avenues for future work in studying LLMs and content moderation.
Introduction
Content Warning: When necessary for clarity, this paper directly quotes user content that contains offensive/hateful speech, profanity, and other potentially triggering content.
Since the release of OpenAI’s ChatGPT in November 2022, publicly available large language models (LLMs) have exploded in popularity and achieved widespread recognition. Such models, like OpenAI’s GPT-suite and Google’s Bard, offer new and accessible ways for end-users to interface with machine learning and AI as well as provide developers with a new toolkit with which to build AI-powered technologies. Indeed, developers are already using LLMs for a variety of text-based tasks: Microsoft incorporated GPT-4 into Bing to enhance web search functionality, [^1] https://www.theverge.com/2023/3/15/23641683/microsoft-bing-ai-gpt-4-chatbot-available-no-waitlist Expedia released a GPT-3.5 powered chatbot to help end-users create travel plans, [^2] https://mashable.com/article/expedia-chatgpt-in-app-travel-booking and Slack incorporated a variety of LLMs to provide summaries of meetings and increase end-user productivity. [^3] https://slack.com/blog/news/introducing-slack-gpt
One use case for LLMs that has achieved significant attention is automated content moderation. Although introducing AI to content moderation is not a new idea (Gillespie 2020), recent research has started investigating how best to incorporate LLMs into content moderation pipelines (Franco, Gaggi, and Palazzi 2023) and industry players alike are expressing enthusiasm for LLM-powered content moderation (Gilardi, Alizadeh, and Kubli 2023). OpenAI recently published an article outlining how GPT-4 could be used to build and evaluate content policies with test examples [^4] https://openai.com/blog/using-gpt-4-for-content-moderation to minimize time required to develop platform policy and reduce potential emotional burden for human reviewers. Given this recent attention, it behooves the research community to investigate the role that LLMs might play in automated content moderation, both in where LLMs excel and where there are opportunities for improvement.
In this paper, we evaluate the effectiveness of several publicly available LLMs on two common content moderation tasks: rule-based community moderation and toxic content detection. For each task, we develop a test harness that combines real-world datasets with careful prompt design and evaluate each model against available ground truth. We focus on OpenAI’s GPT models (3, 3.5, 4) as exemplars of deployed language models. For the rule-based community moderation task, we build 95 LLM-powered moderation engines that are each prompted with community guidelines sourced from 95 subcommunities on Reddit. We prompt each LLM-engine to simulate moderation decisions on comments curated by Chandrasekharan et al. (2018) and compare how well LLM decisions align with human-moderator decisions. For toxic content detection, we construct a balanced dataset of 10K comments sourced by Kumar et al. (2021) and compare GPT-3, GPT-3.5, and GPT-4 performance with ground truth and other popular state-of-the-art commercially available baselines.
We observe that LLMs can be effective at conducting rule-based moderation, achieving an accuracy of 64% and precision of 83% for the median subreddit. For some communities, GPT-3.5 achieves near human-moderator levels of performance: for example, the model achieves an accuracy of 82% and a precision of 95% on comments from r/movies. We investigate what rules GPT-3.5 evokes when simulating moderation decisions, and find that the model is more likely to moderate based on restrictive rules (e.g., those that restrict behavior) compared to prescriptive rules or formatting based rules, suggesting that framing policies as restrictions may be an effective strategy for moderators and policymakers seeking to use LLMs for moderation tasks. We also highlight potential pitfalls with using LLMs for moderation: we observe stark differences in performance when underlying models change over time, which can lead to unpredictable results.
With toxicity detection, we observe that modern LLMs like GPT-3, GPT-3.5, and GPT-4 outperform existing commercial toxicity detection classifiers, with GPT-3.5 achieving the most balanced performance across all tested models (Accuracy of 0.74, F1 of 0.73). Interestingly, we find only marginal improvement in toxicity detection performance when comparing GPT-3, GPT-3.5, and GPT-4, suggesting that increased LLM-size may not always lead to increased benefits in a content moderation context. Still, LLMs are not perfect, and exhibit some significant issues: models often miss implicit toxicity, which has been a topic of recent study in toxicity detection (ElSherief et al. 2021), and, like existing classification-based approaches, at times cannot make accurate decisions without appropriate conversational context.
Our results paint a tempered but optimistic view of introducing LLMs into content moderation contexts. LLMs provide some improvements over existing baselines but will still require human review and careful guardrails before widespread adoption. We conclude with a discussion of our results and chart a pathway for future research in this space. We will make our code and data available at publication time. We hope our results will serve useful to researchers and practitioners investigating how best to incorporate LLMs into content moderation pipelines.
Background and Related Work
Recent research has investigated the role that machine learning and AI-assisted tools can play in tackling content moderation problems. In this section, we provide the necessary background and detail the relevant research we build on to conduct our analyses.
AI-Assisted Content Moderation
As machine learning and AI tools have become ubiquitous, recent research has investigated the role that such tools might play in content moderation. One class of work has examined the tensions and affordances of adding machine learning to content moderation workflows, demonstrating that while automation can enable new ways of tackling content moderation (e.g., through delegation, scale, third-party tooling), in many cases, human moderators still want to make decisions that fall into gray areas (Kuo, Hernani, and Grossklags 2023; Ma and Kou 2023; Atreja et al. 2023; Lai et al. 2022; Jiang et al. 2023; Han et al. 2023). Other work has focused on the design space of adding AI systems to content moderation tasks, particularly through the lens of transparency in moderation decisions (Ma and Kou 2023; Choi and Lee 2023; Jhaver, Bruckman, and Gilbert 2019).
These studies demonstrate how transparency through explanation often leads to prosocial outcomes in online communities, but challenges with scale often preclude such explanations from widespread adoption. Finally, recent systems-focused work examined how to build fully-fledged AI systems that support existing content moderation needs. Chandrasekharan et al. proposed Crossmod (2019), a system that surfaced potential candidates for moderation to community moderators on Reddit based on outcomes in other subcommunities, and showed significant alignment between the AI system and human-moderator decisions. Our work builds on this prior research by examining how well LLMs fare at both enforcing written community guidelines as compared to human moderators, and examining if LLMs can provide useful explanations of its moderation outcomes.
Detecting Toxic Content
Google Jigsaw’s Perspective API (Google Jigsaw 2018) has emerged as the de facto classifer for detecting toxic content online, and it has been used in dozens of research papers to identify toxic content in isolation (Xia et al. 2020; Kumar et al. 2021; Lambert, Rajagopal, and Chandrasekharan 2022; Kumar et al. 2023). There has been significant additional research on improving detection beyond the Perspective API. Hanley et al., for example, demonstrate how a DeBERTa classifier trained with contrastive learning marginally outperforms the Perspective API (2023), and, in a forthcoming paper, He et al. show how prompting local LLMs (e.g., T5, GPT2-M and GPT2-L) can improve performance for a variety of toxicity related tasks (2024). While multiple studies have considered how older LLMs (i.e., GPT-3) can be used to identify hate speech (Chiu, Collins, and Alexander 2021; Schick, Udupa, and Schütze 2021), no work has studied modern LLMs (GPT-3.5, GPT-4) or compared how performance on toxicity detection has changed over time.
LLMs and Rule-Based Content Moderation
In this section, we study how LLMs can aid in rule-based moderation settings. We prompt OpenAI’s GPT-3.5 with lists of rules collected from 95 subreddits and ask the model to simulate moderation decisions for comments sourced from each subreddit. We evaluate the effectiveness of LLMs by comparing simulated decisions to ground-truth decisions made by human moderators.
Methods
Collecting ground-truth moderation decisions
To compare how well LLMs align with human-provided moderation decisions, we leverage a dataset created by Chandrasekharan et al. (2018) that contains comments that were moderated from 100 popular subreddits between 2016–2017. Since their study was conducted, 5 subreddits have gone offline or been removed from the platform, so our study focuses on the remaining 95 subreddits. For each subreddit, we sample 500 comments that were removed by moderators as well as 500 comments that were not removed by moderators during the same time period, thus creating a balanced dataset of moderation decisions for each subreddit. These datasets serve as our baseline for rule-based evaluation.
Extracting Subreddit metadata
We first extract the description of each subreddit using the Reddit API. We then extract the moderation rules of that community by using the Reddit rules API, [^5] https://www.reddit.com/dev/api/#GET_r_subreddit_about_rules which provides a list of rule texts and whether the rule applies to a post, comment, or all content posted to the subreddit. Given we are only investigating comment moderation in this study, we filter any rules that do not apply to comments. We extract the full rule text for each rule and build a numbered list of rules that we use to prompt each model.
Prompt design
We construct a new zero-shot prompted LLM-engine for each subreddit we investigate. We begin by prompting the LLM with the description of the subreddit as well as the constructed rules string for each subcommunity. The full initial prompt text is:
You are a helpful content moderation
assistant for the online subreddit
<subreddit>. The subreddit has the
following description: <description>.
Here are the rules for the subreddit:
<rule_texts>
We then ask the model to evaluate whether it would moderate new unseen comments (yes/no), the rule text and rule numbers that the comment violates, and an explanation for the decision. Similarly to the toxic content detection task, we ask the model to return a JSON object for downstream analysis. The full prompt text is:
Consider the following comment:
<example_comment>
Return a JSON object with four fields,
"would_moderate," that is
either "yes" or "no" depending
on if you would remove this comment
from the subreddit, "rules" which are
the text of the rules being violated,
"rule_nums" which are a comma-separated
list of rules being violated, "rating"
which is a score from 1--5 on how
violative the comment is.
Evaluation strategy
We compare ground-truth moderation decisions with decisions made by each subreddit-specific tuned LLM on a number of binary performance metrics. For this experiment, we solely rely on OpenAI’s GPT3.5 for evaluation as it balances at-scale measurements (i.e., 100,000 queries in total) with cost. To understand why the LLM makes mistakes, we also perform a case study on mistakes created in the r/worldnews subreddit and provide a qualitative analysis of existing gaps in rule-based moderation.
Figure 1: Rule-based moderation aggregate performance by Subreddit—We show a CDF of accuracy, precision, and recall across the 95 subreddits we study. There is considerable spread in performance across subreddits, with GPT-3.5 performing worse than random chance on some and performing near identically to human moderators for others.
Aggregate Performance
To begin, we measure the binary performance of each LLM-engine prompted with each subreddit’s rules compared to human moderators. GPT-3.5 achieves a median accuracy of 63.7%, a median precision of 83%, and a median recall of 39.8% across all subreddits in our dataset (Figure 1). For each metric, there is considerable spread: GPT-3.5 achieves high accuracy for subreddits like r/movies (81.5%) and r/OldSchoolCool (76.3%), with each of these also exhibiting high precision (94% and 82%, respectively). Such performance is not uniform across all subreddits. GPT-3.5 performs worse than a coin toss when simulating decisions from subreddits like r/askscience (Accuracy 38.5%) and r/AskHistorians (Accuracy 41.7%)—suggesting that LLMs are not always a useful community moderation-aid. Notably, both of these subreddits are centered around discussion with experts, and moderation decisions in these subcommunities often require significant context, which is a limitation of our evaluation setup (Gilbert 2020).
Figure 2: Threshold analysis for rule-based moderation—We show the effect of introducing LLM-created thresholds to rule-based moderation tasks across Accuracy, Precision, Recall, and F1. Thresholds in this context are not particularly sensitive, with thresholds from 1–3 achieving near identical results. As such, rating-based thresholding is likely a less effective strategy to enable more control over rule-based moderation than a simple binarized decision.
We next evaluate if thresholding (i.e., asking the model to provide a rating, and using that rating to binarize decisions) can be used to improve or tune rule-based performance, as has been useful in other content moderation contexts (Kumar et al. 2021). Figure 2 shows the binary metrics we use in this paper at each threshold from 0–4. Interestingly, we find that overall performance is not sensitive to thresholding—GPT-3.5 performed near identically for the middle three thresholds, with an accuracy of approximately 62.6% and an F1 of 0.52–0.54. At the highest threshold (t=4), precision greatly increased (0.94), but recall significantly diminished (0.06). Our results suggest that thresholding is not a useful strategy for rule-based moderation than in other contexts.
Errors in Rule-Based Moderation
When GPT-3.5 misaligns with human moderators, it is much more likely to create a false negative (86.9% of all errors) than a false positive (13.1% of all errors). The error distribution for the median subreddit is similar, with 12.1% of errors being false positives and 87.8% of errors being false negatives. Similarly to our prior results, there is significant distributional spread and several subcommunities buck this trend: for example, when moderating the subreddit r/NSFW_GIF, an adult community for sharing not-safe-for-work imagery, GPT-3.5 creates 68% false positives and 31% false negatives, often due to missing context or the model erroneously flagging sexual content, which is desired in this community. These results suggest that while GPT-3.5 favors precision over recall in aggregate, these results do not uniformly impact all subcommunities and depend on content, context, and community norms.
GPT-3.5/030123 GPT-3.5/061323
Accuracy
68% 63%
Precision
75% 83%
Recall
63% 40%
F1
68% 53%
Table 1: Longitudinal LLM performance—We show the median performance of GPT-3.5 across the first and second released models (from March 1, 2023 to June 13, 2023). We find that GPT-3.5 produces drastically different results, with new models significantly preferring precision over recall. Our results suggest that human-moderators may need to stay significantly in the loop as models grow and evolve.
Stability of LLM Performance Over Time
Given LLMs like GPT-3.5 are constantly undergoing changes due to model updates and deployment changes, we also sought to understand how stable model performance is at rule-based moderation tasks over time. Table 1 shows the median performance across all subreddits in our dataset from the first released GPT-3.5 model (released on March 1, 2023) and the second released model (released on June 13, 2023). Performance significantly changes between the two models, despite there being no other differences in prompts or analyzed comments. F1 performance decreased from 0.68 to 0.53, primarily due to a decrease in median recall: which similarly decreased from 0.63 to 0.40. Median precision, however, increased from 0.75 to 0.83—suggesting that the newer model favored reducing false positives over reducing false negatives. Although it is unclear why the underlying GPT-3.5 model shifted in this direction, our results highlight that content moderators wishing to incorporate LLMs into their moderation workflows will need to consider model stability as an additional factor when deploying these tools.
LLMs and Rule Enforcement
LLM decisions are not only governed by the content, but also by the types of rules that are available when making moderation decisions. As part of our prompt design, we ask the LLM to express which rules are violated when it flags a comment for moderation. We use these rules to study what kinds of rules LLMs may be better at expressing over other rules, which can inform how LLMs can be better used in practice by community moderators. To increase our descriptive power for each rule, we replicate the work of Fiesler et al. (2018), who created a model and taxonomy for classifying Reddit rules into several broad categories. We classify each subreddit rule into three categories: restrictive rules (i.e., those that restrict types of behavior), prescriptive rules (i.e., those that tell someone what to do), and format related rules. We detail our model replication process in the appendix.
GPT-3.5 expresses a significant fraction of rules: of the 562 rules in our dataset, GPT-3.5 expressed 467 (83.1%) when making moderation decisions. The median subreddit has 88% of its rules exercised by the LLM, with 49% of subreddits having 100% of their rules expressed by the LLM. Still, there is significant distributional spread, with some subreddits having only a small fraction of their rules expressed. For example, r/AskTrumpSupporters, a subreddit for asking questions to those that supported Donald Trump’s presidency, only has 3 of its 7 rules expressed.
Rule Type
% Overall Rules % Subreddits % Expressed
Restrictive
460 (82.4%) 90 (95%) 398 (87%)
Prescriptive
319 (57.1%) 86 (90.1%) 246 (77%)
Format
46 (8.2%) 28 (29%) 33 (71%)
Table 2: Rule type expression—We show the distribution of rule types across rules and subreddits, as well as how frequently each type was expressed by the LLM when making moderation decisions. LLMs are most likely to moderate via restrictive rules (87%) compared to prescriptive rules (77%) and format-driven rules (71%).
Despite high rule coverage overall, GPT-3.5 does not express all rule types uniformly. Table 2 shows the distribution of rule-types in our dataset, across all rules, the fraction of subreddits that contain at least one rule of each type, and the fraction of rules expressed for each rule type. GPT-3.5 was much more likely to enforce restrictive rules (87%) than prescriptive rules (77%) and format-related rules (71%); in both cases, the LLM is statistically significantly less likely to express either rule type when compared to a restrictive rule per a 2-tailed proportions test ($p<0.01$). Our results suggest that GPT-3.5 are more likely to take action for restrictive rules, suggesting that policymakers wishing to use LLMs for rule-enforcement might need to consider more restrictive framings of rules versus prescriptive framings.
Finally, we investigate if errors made by the LLM in moderation decisions are imbalanced across rule types. Understanding this can illuminate whether LLMs might be better suited to moderate certain types of rules versus other rules. Given that the LLM only expresses a rule if it chooses to moderate the comment, we measure the error rate for each rule type by the fraction of comments for each rule type that the LLM made a false positive. Surprisingly, we find little evidence that rule types impact performance: we find almost no difference in error rates for restrictive rules (18.5% errors), prescriptive rules (15.9% errors) or formatting rules (16.9% errors), highlighting that GPT-3.5 erroneously flags content for moderation across all rule types evenly.
Case Study: r/worldnews
Reason
% Comments
Mistaken Hate
47%
Correct
23%
Context
20%
Sarcasm
16%
Mistaken Generalization
1%
[…]
Reason
% Comments
Context
54%
Missing Attack
30%
Sarcasm
5%
Meme
3%
[…]
Table 3: Explaining errors in moderating r/worldnews—We show the most frequent explanations for errors generated by GPT-3.5 when moderating r/worldnews. False positives are largely due to incorrectly flagging a comment as hate speech, especially when identity language appears (47%). False negatives occur most often when there is missing additional context that needs to be included to make a moderation decision (54%).
We complement our quantitative analyses with a qualitative analysis on errant decisions made by GPT-3.5 on the subreddit r/worldnews, which is a subreddit dedicated to major news around the world. We chose this particular subreddit as it is popular (33M subscribers), it well-balances false positives (33%) and false negatives (67%) to analyze, and tends to more general discussion versus niche subcommunity interests. r/worldnews has only 3 rules that apply to comments, two of which are related to personal attacks and hate speech, and one which is related to meme-content. We manually investigate the 244 comments that produced either a false negative or false positive, iteratively develop a codebook for explaining each error, and manually code each comment with an explanation. Table 3 shows a distribution of common explanations for false positives and false negatives generated by model. We discuss each in detail below:
False positives
The primary reason for false positives is due to GPT-3.5 mistakenly identifying a comment as hate speech (47%). This often occurs when identity language is prevalent in the comment and the sentiment is not overtly clear. For example, the comment “Are you aware that there are other businessmen in the world besides your Jew Boogeyman?” was flagged as violative of hate speech rules, despite the fact that it is attempting to rebuff the original commenter and defend Jewish people from unwanted stereotypes. Given that hate speech and hateful language is part of the Reddiquette, [^6] https://support.reddithelp.com/hc/en-us/articles/205926439-Reddiquette which governs platform-wide norms and behaviors, errors produced by LLMs like these will have a wide impact on all communities wishing to curb and moderate hate speech. Interestingly, we also observe 23% of “false positives” where we agree with the model decision over the human-rater. For example, the comment “Lol you’re a fucking idiot” was not flagged by human moderators, but clearly violates the second rule of r/worldnews. The existence of discrepancies like this is encouraging, as it suggests LLMs can also potentially be used as a double-check aid for human moderators who may also make mistakes when moderating content.
False negatives
The primary reason for false negatives is due to missing context around the conversation (54%). This is a well-known issue in deploying algorithmic content moderation (Gorwa, Binns, and Katzenbach 2020), and not necessarily surprising that LLMs suffer from the same problems when tackling comments in isolation. 30% of false negatives, however, were due to the LLM missing a personal or identity attack. In some cases, these attacks were indirect, for example, one comment begins with “She’ll never do time, don’t be a retard…” has an attack on the original commenter, but it is hard to immediately discern. Other comments, however, are direct attacks, and it is not immediately clear why the model misses these, for example “Not sure what that has to do with free speech you moron. Are you equating murder to speech?” Such errors suggest that even with LLMs, human-review is still important to address algorithmic mistakes.
LLMs and Toxicity Detection
In this section, we investigate how well commodity LLMs perform at toxicity detection when compared to commercially available baselines and identify opportunities to improve LLM performance for toxicity detection.
Methods
Dataset
We leverage a dataset of toxicity annotations curated by Kumar et al. (2021) that consists of 107K comments sourced from three social platforms: Reddit, Twitter, and 4chan. Each comment was labeled for toxicity by five raters on a 5-point Likert scale from “not at all toxic” to “very toxic.” We collapse this score into a binary outcome based on the median toxicity rating. If the median rating was equal to or above “moderately toxic” we denote the comment as toxic, and if the rating was below this threshold we denote the comment as non-toxic. Due to cost limitations, we do not evaluate each LLM on the entire dataset. Rather, we sample the aggregate dataset to build a balanced dataset of toxicity annotations that contains 5000 examples of toxic content and 5000 examples of nontoxic content. We will release this golden dataset at publication time.
Prompt design
We first prompt each LLM with the definition of toxic content used by Google Jigsaw, which is “a rude, disrespectful, or unreasonable comment that is likely to make someone leave a discussion”  (Google Jigsaw 2018). We then ask the model to rate the toxicity of unseen comments on a scale from 1–10, as well as to provide an explanation for the rating in natural language. For ease of downstream analysis, we prompted the LLM to always return results as a JSON object with consistent indexing. The full prompt text is:
Consider the following comment:
<example_comment>
Return a JSON object with a two fields,
"rating," that is a score from 1 to 5
depending on how toxic the comment is
based on our definition, and "explanation"
which contains a reason for the rating.
Evaluation strategy
We evaluate three state-of-the-art LLMs for this task—OpenAI’s GPT3, OpenAI’s GPT3.5 (i.e., ChatGPT), and OpenAI’s GPT4. We chose these as they represent recent, publicly available LLM offerings on the market and covers a range of model sizes and training strategies. To compare across a SOTA commodity baseline, we also label each comment using the Google Jigsaw Perspective API (Google Jigsaw 2018) using their TOXICITY and SEVERE_TOXICITY classifiers at varying thresholds.
Model Threshold Accuracy Precision Recall F1
Perspective API TOXICITY 0.5 0.66 0.68 0.61 0.64
Perspective API TOXICITY 0.7 0.59 0.76 0.27 0.4
Perspective API TOXICITY 0.9 0.52 0.92 0.05 0.1
Perspective API SEVERE_TOXICITY 0.5 0.51 0.96 0.02 0.05
OpenAI GPT-3 0.71 0.71 0.73 0.72
OpenAI GPT-3.5 0.74 0.76 0.7 0.73
OpenAI GPT-4 0.74 0.79 0.65 0.71
Table 4: Toxicity detection aggregate performance—We show aggregate performance from the Perspective API’s TOXICITY and SEVERE_TOXICITY models at varying thresholds (0.5, 0,7, and 0.9) as well as OpenAI’s GPT-3, GPT-3.5, and GPT-4. Across almost every metric, the large language models we tested outperformed the Perspective API at toxicity detection. Surprisingly, we saw only marginal improvements at this task when comparing GPT-3 to GPT-3.5 and GPT-4, highlighting a potential limit to what LLMs might be able accomplish in toxicity detection.
Aggregate Performance
LLMs outperform the Perspective API at toxic content detection across almost all metrics, chiefly in accuracy and in balancing the precision-recall tradeoff (Table 4). OpenAI’s GPT models achieve an accuracy of 0.71–0.74, compared to a maximum accuracy of 0.66 from the Perspective API’s TOXICITY classifier at a threshold of 0.5. The GPT models well-balance precision and recall, achieving F1 scores of 0.71–0.73, compared to a maximum F1 of 0.64 from Perspective. Prior work has noted that Perspective tends to favor precision over recall (Kumar et al. 2021), and we observe a similar trend here: at a high enough threshold with a stringent model (SEVERE_TOXICITY), the Perspective API achieves near perfect precision (0.96) though at a significant impact to recall (0.02).
Interestingly, there is only marginal improvement on this particular toxicity detection task between GPT-3, GPT-3.5, and GPT-4, despite significant enhancements in model sizes and capabilities from GPT-3 to GPT-4. Although GPT-3.5 and GPT-4 return the highest accuracy and F1 scores (0.73), GPT-3’s is comparable, achieving an F1 of 0.72. However, of the LLMs tested, GPT-4 achieves a significant improvement in precision over GPT-3 (0.79 vs. 0.71), suggesting that larger models may greatly favor reducing false positives over reducing false negatives.
Thresholding LLMs for Toxicity Detection
For many toxic content detection tasks, it is often useful to be able to “threshold” a toxic content classifier towards higher recall or higher precision depending on the use-case. For example, prior research has favored a high precision signal when investigating toxic behaviors online (Kumar et al. 2023) and used a high threshold for detection, whereas other research favored F1 for a balance between precision and recall (Saveski, Roy, and Roy 2021) and used a lower threshold for detection. Although LLMs are not explicitly designed as classifiers, we can simulate a classification-like thresholding by selecting a threshold from the score produced by the LLM (1–10). Although thresholding is not an effective strategy in rule-based moderation, we next investigate whether thresholding LLMs can provide more granularity to end-users in identifying toxic content.
Figure 3: Toxicity detection ROC curves—We show the ROC curves for all toxicity detection systems we evaluate. LLMs from OpenAI outperform the Perspective API in AUC, achieving an AUC of 0.8–0.81 compared to a maximum AUC of 0.76 from the Perspective API.
We observe that varying the thresholds for LLMs does produce an internally consistent precision/recall tradeoff. Figure 3 shows the resultant receiver operating characteristic curve for GPT-3.5, GPT-4, and the Perspective API models. The area under the ROC curve (ROC_AUC) for GPT-3.5 and GPT-4 surpasses the AUC of the Perspective API models (0.80 compared to 0.76), demonstrating that the LLMs are better able to balance precision and recall better than Perspective models can. As such, prompting LLMs to provide a discrete rating for toxicity can be another useful lever to content moderators looking to curb toxic content.
Threshold
GPT-3.5 GPT-4
Prec. Recall F1 Prec Recall F1
2
0.53 0.98 0.69 0.58 0.95 0.72
3
0.61 0.92 0.73 0.64 0.88 0.74
4
0.67 0.87 0.75 0.68 0.83 0.74
5
0.68 0.84 0.75 0.71 0.78 0.75
6
0.7 0.81 0.75 0.74 0.73 0.74
7
0.74 0.74 0.74 0.77 0.68 0.72
8
0.8 0.58 0.67 0.84 0.52 0.64
9
0.88 0.34 0.49 0.93 0.27 0.42
10
0.88 0.34 0.49 0.94 0.25 0.39
Table 5: Thresholding LLMs for toxicity detection—Both GPT-3.5 and GPT-4 produce an internally consistent precision/recall tradeoff when performing toxicity detection, highlighting how such models can serve similar functions to toxicity detection classifiers.
We next investigated how increasing classification thresholds for GPT-3.5 and GPT-4 impacts precision, recall, and F1 performance (Table 5). At the highest thresholds, GPT-3.5 and GPT-4 are able to achieve comparable precision (0.94) to the Perspective API (0.96), but without as a significant hit to recall (0.25 vs. 0.02). GPT-3.5 is not able to achieve similarly high precision as GPT-4 or Perspective’s SEVERE_TOXICITY classifier (max 0.88), however, it does well-balance recall with precision (F1 of 0.49, t=10), highlighting that lower-cost LLMs can be equally effective at detecting toxic content in settings where balancing precision and recall is critical.
Do LLM Toxicity Ratings Align with Human Raters?
Although many use-cases of toxicity classifiers in content moderation settings are binary, the severity of the comment also plays a role in downstream moderation action. Recent research has demonstrated that LLMs are comparable to crowdsourced human-raters at many tasks (Törnberg 2023; Gilardi, Alizadeh, and Kubli 2023), however, this has yet to be tested in a toxicity detection setting.
To measure how well LLM toxicity ratings align with human-raters, we compute a Pearson correlation between the rating provided by the LLM (from 1–10) and the median toxicity score across all five raters for each comment (from 1–5) for all comments labeled by ChatGPT and GPT-4. We find there is a strong correlation (GPT-3.5, $r=0.60$, GPT-4, $r=0.61$) between LLM decisions and human raters, suggesting that LLM ratings can also provide a rough proxy for the severity of toxic comments for end-users. For downstream moderation tasks, nuanced notions of severity of content can serve as an additional signal to enable automated, multilayered content moderation, wherein the severity of the content is taken into consideration when deciding the moderation action to take (OpenWeb 2021).
Tag
% Comments
False Target
47%
Language
42%
Stereotype
4.2%
Sexual
4%
Tone
2%
Context
1%
[…]
Tag
% Comments
Not Explicit
49%
Opinion
22%
Context
9%
Positive
6%
Neutral
5%
Humor
4%
[…]
Table 6: Explaining errors in toxicity detection—We show the most common explanations for errors generated by GPT-4. False positives are largely due to misidentifying targets of harassment and offensive language, whereas false negatives are caused by a failure to identify toxic comments without direct threats and overvaluing end-user opinions.
Where Do LLMs Make Mistakes?
Finally, we investigate what gaps still exist when using LLMs for toxic content detection. To study this problem, we randomly sample 200 comment decision mistakes from GPT-4: 100 of which are false positives (e.g., the LLM erroneously flags a comment as toxic) and 100 of which are false negatives (e.g., the LLM fails to catch a toxic comment.) For each of these comments, we then examine the LLM-provided “explanation” to derive insight into how the LLM is making each erroneous decision. We synthesize each explanation into one or two high-level categories: Table 6 shows the distribution of explanations provided for false positives and false negatives in our sample. We describe each class of errors below:
False positives
The primary reason LLMs erroneously flag benign content is due to erroneously identifying subjects of a comment as targets of harassment (47%). For example, a comment that discussed the coaching of a sports team (“Not the players bro.. Buh [sic] foolish coach”) was flagged as toxic due to the coach finding the comment hurtful, despite the fact that the coach of the team was not a part of the online conversation. Similarly, GPT-4 is at times attuned to poor language (e.g., profanity, slurs) and stereotypes, even in instances of neutral or positive connotation. For example, the comment “Yo n***a wears jean shorts” was not found to be toxic by human raters, but the model flagged this comment for “using derogatory racial slurs.” Prior work has highlighted how toxicity models can inherit unintended social biases (Dixon et al. 2018; Halevy et al. 2021)—such considerations should also apply when leveraging LLMs for content moderation focused tasks.
False negatives
The top reason that LLMs miss toxic content flagged by human raters is that the comment contains no explicit or direct threat or insult to another individual (49%). For example, the comment: “There is hunger and illiteracy in the land. Only in Nigeria.” was not flagged by the model, due to there being no personal attacks. However, the comment still generalizes about the country of Nigeria, and human raters found this comment to be toxic. Similarly, the model is also reticent to flag what it deems to be opinions as toxic (22%), especially when the opinion is couched in a question. For example, the comment “They offered the cheapest goods for a long time in the US. Why do you think that’s not beneficial?” was not flagged as toxic, despite prior work identifying that rhetorical tone can lead to increased toxicity and conversational collapse (Zhang et al. 2018).
Discussion and Conclusion
Our work studied how well LLMs perform on two content-moderation related tasks: rule-based community moderation and toxic content detection. Our results provide insight in where existing LLMs excel and informs what gaps need to be addressed for such tools to have widespread impact in content moderation contexts. We now discuss our results, outline future directions for research in LLM-powered content moderation, and highlight limitations of our approach.
Increasing transparency in LLM decisions
LLM performance can rapidly change over time, as we observed in our rule-based moderation experiments. The root cause of why model changes impact moderation decisions are unclear, but our results align with prior work in studying longitudinal LLM behavior inconsistencies (Chen, Zaharia, and Zou 2023). Content moderators seeking to incorporate LLMs must exercise caution when deploying such tools in practice to maintain the reliability and stability of their moderation decisions. One area of future work is thus to identify tasks that change drastically when models change, to develop better benchmarks for continuous monitoring of LLMs.
Improving rule-based prompting with behavior
Rule-based LLM performance falls below existing AI-assisted content moderation tools. For example, Chandrasekharan et al.’s CrossMod tool (2019), which is developed using the same dataset as the one we leveraged in this paper, is able to achieve an accuracy of 86%, compared to GPT-3.5’s 64%. One potential reason for this is that while LLMs are able to reason in a “forward direction,” by interpreting rules and examining if content falls within those rules, CrossMod is a “reverse direction” tool which starts from actual content removals and learns patterns that can be applied to other communities. An area of future work that might be promising is to see if combining rule-based reasoning with few-shot prompting behavioral examples might improve overall LLM performance without requiring any additional training or fine-tuning of the underlying models.
Performance saturation in toxicity detection
Although LLMs outperform existing commercial toxicity detection tools, we observed that LLM performance only marginally improves at toxicity detection despite the fact that newer model sizes are increasing by orders of magnitude (Bastian 2023). As such, our results paint caution to the notion that increasing model sizes will necessarily lead to downstream improvements in content moderation tasks. One potential area of future work can be in identifying what aspects of toxicity detection remain challenging in spite of using an LLM (e.g., implicit hate or identity language) and potentially offload gray area comments to either more specialized automated pipelines or human review.
Parsing conversational context with LLMs
An ongoing challenge with all automated content moderation tools is in implementing conversational context. In our work, we found missing context to be a prevalent reason for false negatives and false positives in both tasks. Embedding context into toxicity detection has proved challenging in prior work (Pavlopoulos et al. 2020). However, most LLMs are already designed to be conversational in nature, and another potential area for future work might be identifying how best to leverage the conversational interface to simulate both conversation and moderation decisions.
Limitations of our work
This work primarily focuses on text-based content moderation tasks, and specifically on comment moderation on discussion oriented platforms. As such, our results may not extend to other types of moderation, for example, those involving mixed modalities (i.e., images, videos, etc.) or those on social platforms with other affordances (i.e., real-time communication). Furthermore, our evaluation here is done on balanced datasets and thus our comment distributions do not reflect comment distributions found in real-world platforms. We also note that while our results are promising, the LLMs we tested are currently too expensive to immediately impact most content moderation contexts, and future research is needed to build models that can balance performance with cost.
Conclusion
Overall, our results provide a tempered but optimistic view of introducing LLMs into content-moderation contexts. We hope that our work provides the groundwork for researchers and practitioners and charts a pathway forward for more work in studying how LLMs can shape and impact the future of content moderation.
Appendix
Rule Type Classifier Details
In order to add more descriptive labels to each rule in our dataset, we replicate the results from Fiesler et al. (2018). We build three logistic regression classifiers for identifying restrictive rules, prescriptive rules, and format rules, as these types of rules were abundant in the training data. We use their training dataset and closely follow the model presented in their paper, which is a logistic regression classifier based on unigram, bigram, and trigram features extracted from each rule text. We held out 20% of rules as a testing set with a random seed of 42. Ultimately, we were able to achieve comparable results to their paper: the restrictive model achieved an F1 of 0.93, the prescriptive model achieved an F1 of 0.89, and the format model achieved an F1 of 0.79.